OpenStudy (anonymous):

Use DeMoivre's Theorem to write the complex number in trigonometric form. (cosπ/4+isinπ/4)^3

OpenStudy (kropot72):

Simply multiply each of the angles by 3 and remove the exponent 3.

OpenStudy (anonymous):

So its (3cospi/12+isin/12)

OpenStudy (anonymous):

3(cospi/12+isinpi/12)

OpenStudy (kropot72):

No. What is (pi/4) * 3

OpenStudy (kropot72):

\[\frac{\pi}{4}\times 3=\]

OpenStudy (kropot72):

(pi/4) * 3 = (3 * pi)/4

OpenStudy (anonymous):

yes..

OpenStudy (kropot72):

Go ahead and substitute that value as the angles

OpenStudy (kropot72):

Don't forget to remove the ^3 exponent after the last bracket.